OpenStudy (anonymous):

Solve for y: 2x + y =4, 3x -2y = -1

OpenStudy (anonymous):

y=4-2x 3x-2(4-2x)=-1->x=1 y=4-2*1=2

OpenStudy (anonymous):

Would you explain a bit more... how do u do that ... how can it be prove ?

OpenStudy (anonymous):

2x+y=4->y=4-2x then let y=4-2x to3x -2y = -1 to get x=1 then let x=1 into y=4-2x to get y=2

OpenStudy (anonymous):

very well... excellent ... you all guys are amazing...

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

What is this equation ... is it Linear Equation?

OpenStudy (anonymous):

yes. it's Linear Equation